OpenStudy (anonymous):

2 logbase8(x) = logbase8(2) + logbase8(3x + 8). solve this analytically?

OpenStudy (anonymous):

log8 x^2 = log8 (6x+16) x^2 = 6x + 16 x^2 - 6x - 16 = 0 (x+2)(x-8) = 0 x = -2 , 8

OpenStudy (anonymous):

\[\log_{8} x^{2} = \log_{8} (2 \times(3x +8)) \]\[x ^{2} = 6x + 16\] on solving the quadratic eqn. => x = 8,-2

OpenStudy (anonymous):

i put in -2,8 and its wrong? are one of those not right?

OpenStudy (anonymous):

actually -2 isn't right as per the given original equation since log is not defined for negative numbers. so the answer is only 8 if u ask me

OpenStudy (anonymous):

yea thats what i thought and yea just 8 is right